OpenStudy (anonymous):

The step to solve 4x2-13=3

OpenStudy (anonymous):

solve for x^2 and the take the square root

OpenStudy (anonymous):

then

OpenStudy (anonymous):

Transfer 13 to RHS and divide the equation by 4 You get x^2 = 16/4 i.e x^2=4 Take square root on both sides. The answer is x=2 or x=-2

OpenStudy (anonymous):

@su225 good but don't give the answer right away. you may want to read the code of conduct

OpenStudy (anonymous):

k..Sorry

OpenStudy (anonymous):

no worries!